# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ 1 Introduction
Photoluminescent defects, in diamond, hexagonal boron nitride and silicon carbide with optically addressable spins are vital for a broad range of emerging room-temperature quantum technologies. The nitrogen vacancy (NV) center in diamond is the most technologically advanced defect and has been used for many imaging and sensing applications [1, 2], most of which are based on bulk diamond samples. However, diamond is hard and brittle, relatively expensive, and today’s quantum diamond chips are only available up to 4 mm in size. This limits their integration with other materials and their use in large-area or high-throughput (e.g. point-of-care biomedical devices) applications. Hence, diamond nanoparticles (NPs) on solid substrates [3, 4, 5, 6, 7] or integrated into other materials [8, 9, 10, 11, 12] are explored for these applications. However, while fluorescent nanodiamonds (FNDs) containing NV centers offer many potential advantages, their photoluminescence (PL) brightness is generally lower and their spin coherence times shorter than that of NVs in bulk diamond, limiting their sensitivity.
@@
More recently, spin-active defects in hBN have emerged as a promising 2D material platform for nanoscale sensing [13]. Unlike diamond, many as-synthesized hBN materials from NPs to bulk crystals, contain a range of quantum emitters with PL from the violet to the near-infrared spectral region in the same material [14]. However, the PL properties of individual particles can vary greatly [15] and one defect that can be controllably created in hBN [16], the boron vacancy $V_{B}^{-}$ [17, 18], is very dim and has not been observed as a single photon emitter.
@@
Hence, methods for controlling and enhancing FND and hBN PL properties are highly sought after. One approach is to use optical cavities for Purcell-enhanced PL, which has been explored for emitters in diamond [19, 20] and hBN [21, 22]. For diamond, this has been explored for individual FNDs in photonic crystal cavities [23, 24, 25, 26], in metamaterial photonic cavities [27], in open cavities [28, 29], and photonic crystal cavities in bulk diamond [30, 31, 32, 33]. For hBN, this enhance ment approach has been explored in planar Fabry-Pérot–based cavities [34, 35, 36], open cavities [37], plasmonic cavities [38, 39], micronanoresonators [40, 41, 42], and photonic crystal cavities [43, 44, 45, 46, 47]. However, to date, no straightforward and scalable method for the fabrication of thin-film cavities containing nanoscale quantum sensors has been reported. Wafer-scale quantum sensor-doped thin-film cavities can enable, for example, magnetic imaging of currents in microelectronics [48] as illustrated in Fig. 1a.
@@
Here, we demonstrate cavity-enhanced PL of FND and hBN NP quantum sensors embedded in polymer-based thin-film cavities. We report a straightforward and scalable method for the fabrication of quantum sensor-doped thin-film microcavities on the centimeter scale. The PL properties of FNDs and hBN NPs are investigated as a function of the spectral position of the cavity resonance across the visible and near-infrared spectral range. The cavity resonances efficiently modulate the spectral PL peak position of NV centers in FNDs across the NV PL spectrum and lead to an up to 2.9-fold enhancement of the NV PL dacay rate compared to FNDs outside the cavity. The hBN NP PL brightness increases by up to a factor of three and the PL decay rate is enhanced up to 13-fold inside the cavity thin-films. By comparing experimental and theoretical Purcell enhancement factors, we conclude that Purcell enhancement causes the observed modulation of NV PL and that Purcell enhancement alone cannot explain the strong PL brightness and decay enhancement of hBN NP PL. Finally, we demonstrate an up to 4.8 times improved magnetic field sensitivity of 20 nm and 100 nm FNDs in thin-film cavities due to a cavity-enhanced optically detected magnetic resonance (ODMR) contrast.
@@ 2 Results and Discussion
To fabricate the microcavities, we first deposit a fully reflective 100 nm Ag layer on Si via electron beam evaporation physical vapor deposition to create flat, uniform substrates for spin coating. For the FND cavities, $\sim$120 nm FNDs (see Supplementary Information (SI) for details) containing 1 ppm of NV centers were dispersed at a concentration of 0.5 mg/mL in a 1:3 water:1-propanol solution containing polyvinyl pyrrolidone (PVP) at 2.5 wt%. This suspension was spin-coated at speeds from 600 to 2600 rpm to produce FND-doped polymer thin-films ranging in thickness from 107–210 nm. Finally, a semi-transparent Ag layer was deposited via electron beam evaporation to complete the microcavity. For hBN, this fabrication protocol was repeated using as-received hBN particles (Graphene supermarket, 70 nm particle size) dispersed at a concentration of 0.3 mg/mL in a solution containing polymethyl methacrylate (PMMA) in chlorobenzene at 3.0 wt.%. These particles are known to contain a range of photoluminescent spin-active defects that emit from the violet to near-infrared spectral region [14].
Fig. 1d shows a photograph of a $20\times 20$ mm Si-wafer substrate, half of which is coated with a complete FND-doped microcavity (left), while the other half lacks the top Ag mirror for control experiments (right). The light absorption by the resonant cavity mode produces a green tint in the cavity region, which is not present in the no-cavity region. Fig. 1e shows a confocal PL image of the FND microcvity, showing evenly distributed FND particles throughout the field of view. For both FND and hBN microcavities, we fabricated a suite of devices with fundamental optical modes spanning the spectral region from 500–800 nm, covering the spectral emission range of the NV center in FNDs and the hBN emitters. Supplementary Fig. S1a and b show the normal-angle reflectance spectra for the FND and hBN microcavities, respectively, clearly showing reflectance anti-peaks with reflectivity decreases of 60–80% at the predicted resonance mode energies (see Supplementary Fig. S1).
@@
Figure 2: (a,d) Thin-film cavity reflectivity spectra (top, colored traces) and corresponding typical PL spectra of FND (a) and hBN NPs (d) quantum sensors embedded in the cavity (bottom, colored traces). Black traces at the top of a) and b) show the no-cavity reflectivity and no-cavity PL spectra of FNDs (a) and hBN (d) at the bottom. The green bar on the left indicates the excitation wavelength $\lambda_{ex}$= 520 nm. (b,e) Box and whisker plot of $\lambda_{em}$ as a function of $\lambda_{res}$ for all investigated FNDs (b) and hBN NPs (e). The horizontal dash in the box represents the average, the edges of the box the upper and lower quartiles, and the whiskers the minimum and maxiumum values of $\lambda_{em}$. Statistical outliers (1.5 × interquartile range) are in grey. (c,f) Scatter plot of cavity enhacement factor $\eta$ (see main text for details) as a function of $\lambda_{em}$ for individual FNDs (c) and hBN particles (f) , where the color of each dot indicates $\lambda_{res}$ of the cavity.
We then performed the same analysis for hBN NPs. The hBN particles studied here contain a range of carbon-related defects [14], whose spectral PL characteristics vary greatly between individual particles, with emission from 550–800 nm upon 520 nm excitation (see Supplementary Fig. S3 for examples). Fig. 2d shows the PL spectrum of an hBN particle outside the cavity (black trace) compared to spectra of hBN particles in three different cavities (colored traces). Unlike for FNDs, there is no systematic correlation between $\lambda_{res}$ and $\lambda_{em}$. Due to the great variation in hBN PL spectral characteristics, even the identification of a single emission peak position is non-trivial (see Supplementary Fig S4 for details). Fig. 2b shows a box and whisker plot of $\lambda_{em}$ as a function of $\lambda_{res}$ for all investigated hBN particles. It confirms that $\lambda_{em}$ of most hBN particles inside the cavity is between 600–650 nm, irrespective of $\lambda_{res}$, with a large spread of $\lambda_{em}$ values between 570 nm and 700 nm. Interestingly, we nonetheless observe that the PL of more than 90% of hBN particles in cavities is enhanced 2-10 times compared to particles outside the cavity (Fig. 2f). The hBN particles generally appeared slightly larger than FNDs in confocal PL images (see Supplementary Fig. S5) suggesting hBN NP aggregation. Due to their larger size, the aggregates may have been in closer proximity to the top and bottom silver mirror, leading to surface plasmon-related enhancements that are independent of the cavity resonance.
@@ 2.1 Emission Rate Enhancement
The measured PL decay rate $\gamma$ is the sum of radiative $\gamma^{rad}$ and non-radiative decay $\gamma^{nr}$ rates. To identify whether the observed increase in $\gamma$ is the result of an increase in $\gamma^{rad}$, $\gamma^{nr}$ or both, we investigated the total PL brightness $\Phi$ of FNDs, which is proportional to the PL quantum yield $\gamma^{rad}$$\gamma^{rad}+\gamma^{nr}$. We determined $\Phi$ for FNDs inside and outside the cavity as the total area under the PL spectra for each FND. Fig. 3c shows $\Phi_{avg}$ for all devices as a function of $\lambda_{res}$ for FNDs inside and outside the cavity. Markers represent the brightness averaged over 10 FNDs per sample and the error bars the standard deviation between individual FNDs. In half of the devices, FNDs inside and outside the cavity exhibit a similar brightness, with some devices showing a roughly 2-fold decrease ($\lambda_{res}=525$ nm and $\lambda_{res}=700$ nm and above). Importantly, the FND brightness is the same inside and outside the cavity (within the standard deviation) for three devices with $\lambda_{res}$ values of 575–650 nm, for which we see significant decay rate enhancements. Given the observed NV decay rate enhancement of 2.9 observed for $\lambda_{res}=650$ nm and assuming no optical losses in the Ag top mirror and an NV quantum yield of 0.5 [49], this would suggest that both $\gamma^{rad}$ and $\gamma^{nr}$ increase by a factor of 2.9, which would lead to no change in the NV quantum yield and brightness. However, considering light absorption in the Ag mirror is present but difficult to quantify in our cavities, an increase in the radiative rate is likely dominant.
To investigate whether the observed enhancement can be understood in terms of a Purcell enhancement, we determined an experimental Purcell enhancement factor $F_{P}$ [50] and compared it to a theoretical enhancement factor $F_{P}^{th.}$. These are defined as
@@
$F_{P}^{th.}\simeq\frac{6Q}{\pi^{2}n}$ (3)
where $Q$ is the cavity quality factor and $n$ the refractive index of the cavity medium (see Supplementary Information for further details).
Fig. 3e shows the $F_{P}$ and $F_{P}^{th.}$ for all FND cavities as a function of $\lambda_{res}$. Most cavities exhibit a theoretical Purcell enhancement of $F_{P}^{th.}=5\pm 0.5$. The cavity with $\lambda_{res}=650$ nm has a slightly higher $F_{P}^{th.}$of 6.2, coinciding with the highest observed experimental $F_{P}$ of 2.9 in the spectral region of the NV- ZPL. In the spectral region above 700 nm where the NV PL begins to drop off, $F_{P}^{th.}$ slightly increases while $F_{P}$ decreases. In all cases $F_{P}$ is significantly lower than $F_{P}^{th.}$, which is expected since cavity losses, e.g. in the Ag mirrors, are not considered in our model. However, the three devices with resonances from 575–650 nm, for which the highest decay rate enhancement was observed, all show average $F_{P}>1$, suggesting that a Purcell enhancement of the NV radiative rate causes the observed enhancements.
@@
Fig. 3f shows the resulting decay rate enhancement $\Gamma_{avg}$ for all hBN devices as a function of $\lambda_{res}$. We find significant decay rate enhancements for all devices with average values for different devices ranging from 4 to 13 and also strong variations between individual particles within the same device. Fig. 3f shows the corresponding average hBN brightness as a function of $\lambda_{res}$. Unlike for FNDs, we find a systematic increase in PL brightness for hBN particles inside the cavity compared to outside and no systematic dependence on $\lambda_{res}$. We also observed that the hBN NPs inside the cavity showed pronounced fluctuations in PL intensity over time. These fluctuations were also observed in real-time in steady-state widefield measurements and is a behavior not observed in the no-cavity control devices. This blinking may arise from transient switching of defect centers between radiative and non-radiative states [51]. Coupling to the planar metallic microcavity may enhance these effects through resonant interactions, including via Purcell effects and plasmonic coupling, where fluctuations in the charge state of the emitter may be amplified in the cavity, leading to the temporally unstable blinking observed in these measurements. These observations indicate that the cavity not only modifies spectral features but also reveals dynamic emission behaviors that are absent in uncoupled nanoparticles.
@@
Several mechanisms may explain our observations. Both the excitation beam and the hBN PL can be enhanced by cavity resonances. Since we don’t observe a clear dependence of emission enhancement on lambda res, it is unlikely that cavity resonances are the main cause for the enhancement. The close proximity of the hBN particles to the top and bottom silver mirror may cause surface plasmon-related enhancements. Direct electrical contact between hBN particles and one of the mirrors can also lead to a fast light-induced redistribution of charges that may explain the fast PL decay. Lastly, a local change in cavity geometry due to the formation of hBN aggregates could alter the cavity properties around particles and may focus light onto hBN particles independent of the polymer layer thickness. Future studies will investigate the origin of the observed PL modulation.

### v2

@@ 1 Introduction
Photoluminescent defects, in diamond, hexagonal boron nitride and silicon carbide with optically addressable spins are vital for a broad range of emerging room-temperature quantum technologies. The nitrogen vacancy (NV) center in diamond is the most technologically advanced defect and has been used for many imaging and sensing applications Rovny et al. [2024], Aslam et al. [2023], most of which are based on bulk diamond samples. However, diamond is hard and brittle, relatively expensive, and today’s quantum diamond chips are only available up to 4 mm in size. This limits their integration with other materials and their use in large-area or high-throughput (e.g. point-of-care biomedical devices) applications. Hence, diamond nanoparticles (NPs) on solid substrates Heffernan et al. [2017], Kianinia et al. [2016], Shulevitz et al. [2022], Foy et al. [2020], Chea et al. [2025] or integrated into other materials Andrich et al. [2018], Styles et al. [2024], Guarino et al. [2020], Khalid et al. [2020], Price et al. [2019] are explored for these applications. However, while fluorescent nanodiamonds (FNDs) containing NV centers offer many potential advantages, their photoluminescence (PL) brightness is generally lower and their spin coherence times shorter than that of NVs in bulk diamond, limiting their sensitivity.
@@
More recently, spin-active defects in hBN have emerged as a promising 2D material platform for nanoscale sensing Aharonovich et al. [2022]. Unlike diamond, many as-synthesized hBN materials from NPs to bulk crystals, contain a range of quantum emitters with PL from the violet to the near-infrared spectral region in the same material Singh et al. [2025]. However, the PL properties of individual particles can vary greatly Xu et al. [2018] and one defect that can be controllably created in hBN Choi et al. [2016], the boron vacancy $V_{B}^{-}$ Tran et al. [2016], Gottscholl et al. [2020], is very dim and has not been observed as a single photon emitter.
@@
Hence, methods for controlling and enhancing FND and hBN PL properties are highly sought after. One approach is to use optical cavities for Purcell-enhanced PL, which has been explored for emitters in diamond Janitz et al. [2020], Katsumi et al. [2025] and hBN Jang et al. [2023], Wang et al. [2025]. For diamond, this has been explored for individual FNDs in photonic crystal cavities Wolters et al. [2010], Tomljenovic-Hanic et al. [2011], Fehler et al. [2019], Schrinner et al. [2020], in metamaterial photonic cavities Bar-David et al. [2023], in open cavities Kaupp et al. [2016], Jeske et al. [2017], and photonic crystal cavities in bulk diamond Liang et al. [2012], Riedel et al. [2017], Schröder et al. [2014], Berghaus et al. [2025]. For hBN, this enhance ment approach has been explored in planar Fabry-Pérot–based cavities Vogl et al. [2019], Zeng et al. [2023], Scheuer et al. [2023], open cavities Häußler et al. [2021], plasmonic cavities Mendelson et al. [2021], Genc et al. [2025], micronanoresonators Proscia et al. [2020], Parto et al. [2022], Sakib et al. [2024], and photonic crystal cavities Fröch et al. [2022], Nonahal et al. [2023], Qian et al. [2022], Fröch et al. [2020], Kim et al. [2018]. However, to date, no straightforward and scalable method for the fabrication of thin-film cavities containing nanoscale quantum sensors has been reported. Wafer-scale quantum sensor-doped thin-film cavities can enable, for example, magnetic imaging of currents in microelectronics Turner et al. [2020] as illustrated in Fig. 1a.
@@
Here, we demonstrate cavity-enhanced PL of FND and hBN NP quantum sensors embedded in polymer-based thin-film cavities. We report a straightforward and scalable method for the fabrication of quantum sensor-doped thin-film microcavities on the centimeter scale. The PL properties of FNDs and hBN NPs are investigated as a function of the spectral position of the cavity resonance across the visible and near-infrared spectral range. The cavity resonances efficiently modulate the spectral PL peak position of NV centers in FNDs across the NV PL spectrum and lead to an up to 2.9-fold enhancement of the NV PL decay rate compared to FNDs outside the cavity. The hBN NP PL brightness increases by up to a factor of three and the PL decay rate is enhanced up to 13-fold inside the cavity thin-films. By comparing experimental and theoretical Purcell enhancement factors, we conclude that Purcell enhancement causes the observed modulation of NV PL and that Purcell enhancement alone cannot explain the strong PL brightness and decay enhancement of hBN NP PL. Finally, we demonstrate an up to 4.8 times improved magnetic field sensitivity of 20 nm and 100 nm FNDs in thin-film cavities due to a cavity-enhanced optically detected magnetic resonance (ODMR) contrast.
@@ 2 Results and Discussion
To fabricate the microcavities, we first deposit a fully reflective 100 nm Ag layer on Si via electron beam evaporation physical vapor deposition to create flat, uniform substrates for spin coating. For the FND cavities, $\sim$120 nm FNDs (see Supplementary Information (SI) for details) containing 1 ppm of NV centers were dispersed at a concentration of 0.5 mg/mL in a 1:3 water:1-propanol solution containing polyvinyl pyrrolidone (PVP) at 2.5 wt%. This suspension was spin-coated at speeds from 600 to 2600 rpm to produce FND-doped polymer thin-films ranging in thickness from 107–210 nm. Finally, a semi-transparent Ag layer was deposited via electron beam evaporation to complete the microcavity. For hBN, this fabrication protocol was repeated using as-received hBN particles (Graphene supermarket, 70 nm particle size) dispersed at a concentration of 0.3 mg/mL in a solution containing polymethyl methacrylate (PMMA) in chlorobenzene at 3.0 wt.%. These particles are known to contain a range of photoluminescent spin-active defects that emit from the violet to near-infrared spectral region Singh et al. [2025].
Fig. 1d shows a photograph of a $20\times 20$ mm Si-wafer substrate, half of which is coated with a complete FND-doped microcavity (left), while the other half lacks the top Ag mirror for control experiments (right). The light absorption by the resonant cavity mode produces a green tint in the cavity region, which is not present in the no-cavity region. Fig. 1e shows a confocal PL image of the FND microcavity, showing evenly distributed FND particles throughout the field of view. For both FND and hBN microcavities, we fabricated a suite of devices with fundamental optical modes spanning the spectral region from 500–800 nm, covering the spectral emission range of the NV center in FNDs and the hBN emitters. Supplementary Fig. S1a and b show the normal-angle reflectance spectra for the FND and hBN microcavities, respectively, clearly showing reflectance anti-peaks with reflectivity decreases of 60–80% at the predicted resonance mode energies (see Supplementary Fig. S1).
@@
Figure 2: (a,d) Thin-film cavity reflectivity spectra (top, colored traces) and corresponding typical PL spectra of FND (a) and hBN NPs (d) quantum sensors embedded in the cavity (bottom, colored traces). Black traces at the top of a) and b) show the no-cavity reflectivity and no-cavity PL spectra of FNDs (a) and hBN (d) at the bottom. The green bar on the left indicates the excitation wavelength $\lambda_{ex}$= 520 nm. (b,e) Box and whisker plot of $\lambda_{em}$ as a function of $\lambda_{res}$ for all investigated FNDs (b) and hBN NPs (e). The horizontal dash in the box represents the average, the edges of the box the upper and lower quartiles, and the whiskers the minimum and maxiumum values of $\lambda_{em}$. Statistical outliers (1.5 × interquartile range) are in grey. (c,f) Scatter plot of cavity enhancement factor $\eta$ (see main text for details) as a function of $\lambda_{em}$ for individual FNDs (c) and hBN particles (f) , where the color of each dot indicates $\lambda_{res}$ of the cavity.
We then performed the same analysis for hBN NPs. The hBN particles studied here contain a range of carbon-related defects Singh et al. [2025], whose spectral PL characteristics vary greatly between individual particles, with emission from 550–800 nm upon 520 nm excitation (see Supplementary Fig. S3 for examples). Fig. 2d shows the PL spectrum of an hBN particle outside the cavity (black trace) compared to spectra of hBN particles in three different cavities (colored traces). Unlike for FNDs, there is no systematic correlation between $\lambda_{res}$ and $\lambda_{em}$. Due to the great variation in hBN PL spectral characteristics, even the identification of a single emission peak position is non-trivial (see Supplementary Fig S4 for details). Fig. 2b shows a box and whisker plot of $\lambda_{em}$ as a function of $\lambda_{res}$ for all investigated hBN particles. It confirms that $\lambda_{em}$ of most hBN particles inside the cavity is between 600–650 nm, irrespective of $\lambda_{res}$, with a large spread of $\lambda_{em}$ values between 570 nm and 700 nm. Interestingly, we nonetheless observe that the PL of more than 90% of hBN particles in cavities is enhanced 2-10 times compared to particles outside the cavity (Fig. 2f). The hBN particles generally appeared slightly larger than FNDs in confocal PL images (see Supplementary Fig. S5) suggesting hBN NP aggregation. Due to their larger size, the aggregates may have been in closer proximity to the top and bottom silver mirror, leading to surface plasmon-related enhancements that are independent of the cavity resonance.
@@ 2.1 Emission Rate Enhancement
The measured PL decay rate $\gamma$ is the sum of radiative $\gamma^{rad}$ and non-radiative decay $\gamma^{nr}$ rates. To identify whether the observed increase in $\gamma$ is the result of an increase in $\gamma^{rad}$, $\gamma^{nr}$ or both, we investigated the total PL brightness $\Phi$ of FNDs, which is proportional to the PL quantum yield $\gamma^{rad}$$\gamma^{rad}+\gamma^{nr}$. We determined $\Phi$ for FNDs inside and outside the cavity as the total area under the PL spectra for each FND. Fig. 3c shows $\Phi_{avg}$ for all devices as a function of $\lambda_{res}$ for FNDs inside and outside the cavity. Markers represent the brightness averaged over 10 FNDs per sample and the error bars the standard deviation between individual FNDs. In half of the devices, FNDs inside and outside the cavity exhibit a similar brightness, with some devices showing a roughly 2-fold decrease ($\lambda_{res}=525$ nm and $\lambda_{res}=700$ nm and above). Importantly, the FND brightness is the same inside and outside the cavity (within the standard deviation) for three devices with $\lambda_{res}$ values of 575–650 nm, for which we see significant decay rate enhancements. Given the observed NV decay rate enhancement of 2.9 observed for $\lambda_{res}=650$ nm and assuming no optical losses in the Ag top mirror and an NV quantum yield of 0.5 Plakhotnik and Aman [2018], this would suggest that both $\gamma^{rad}$ and $\gamma^{nr}$ increase by a factor of 2.9, which would lead to no change in the NV quantum yield and brightness. However, considering light absorption in the Ag mirror is present but difficult to quantify in our cavities, an increase in the radiative rate is likely dominant.
To investigate whether the observed enhancement can be understood in terms of a Purcell enhancement, we determined an experimental Purcell enhancement factor $F_{P}$ Englund et al. [2010] and compared it to a theoretical enhancement factor $F_{P}^{th.}$. These are defined as
@@
$F_{P}^{th.}\simeq\frac{6Q}{\pi^{2}}$ (3)
where $Q$ is the cavity quality factor (see Supplementary Information for further details).
Fig. 3e shows the $F_{P}$ and $F_{P}^{th.}$ for all FND cavities as a function of $\lambda_{res}$. Most cavities exhibit a theoretical Purcell enhancement of $F_{P}^{th.}=7.5\pm 1.0$. The cavity with $\lambda_{res}=650$ nm has a slightly higher $F_{P}^{th.}$ of 9.7, coinciding with the highest observed experimental $F_{P}$ of 2.9 in the spectral region of the NV- ZPL. In the spectral region above 700 nm where the NV PL begins to drop off, $F_{P}^{th.}$ slightly increases while $F_{P}$ decreases. In all cases $F_{P}$ is significantly lower than $F_{P}^{th.}$, which is expected since cavity losses, e.g. in the Ag mirrors, are not considered in our model. However, the three devices with resonances from 575–650 nm, for which the highest decay rate enhancement was observed, all show average $F_{P}>1$, suggesting that a Purcell enhancement of the NV radiative rate causes the observed enhancements.
@@
Fig. 3f shows the resulting decay rate enhancement $\Gamma_{avg}$ for all hBN devices as a function of $\lambda_{res}$. We find significant decay rate enhancements for all devices with average values for different devices ranging from 4 to 13 and also strong variations between individual particles within the same device. Fig. 3f shows the corresponding average hBN brightness as a function of $\lambda_{res}$. Unlike for FNDs, we find a systematic increase in PL brightness for hBN particles inside the cavity compared to outside and no systematic dependence on $\lambda_{res}$. We also observed that the hBN NPs inside the cavity showed pronounced fluctuations in PL intensity over time. These fluctuations were also observed in real-time in steady-state widefield measurements and is a behavior not observed in the no-cavity control devices. This blinking may arise from transient switching of defect centers between radiative and non-radiative states White et al. [2020]. Coupling to the planar metallic microcavity may enhance these effects through resonant interactions, including via Purcell effects and plasmonic coupling, where fluctuations in the charge state of the emitter may be amplified in the cavity, leading to the temporally unstable blinking observed in these measurements. These observations indicate that the cavity not only modifies spectral features but also reveals dynamic emission behaviors that are absent in uncoupled nanoparticles.
@@
Several mechanisms may explain our observations. Both the excitation beam and the hBN PL can be enhanced by cavity resonances. Since we don’t observe a clear dependence of emission enhancement on $\lambda_{res}$, it is unlikely that cavity resonances are the main cause for the enhancement. The close proximity of the hBN particles to the top and bottom silver mirror may cause surface plasmon-related enhancements. Direct electrical contact between hBN particles and one of the mirrors can also lead to a fast light-induced redistribution of charges that may explain the fast PL decay. Lastly, a local change in cavity geometry due to the formation of hBN aggregates could alter the cavity properties around particles and may focus light onto hBN particles independent of the polymer layer thickness. Future studies will investigate the origin of the observed PL modulation.